OpenStudy (anonymous):

How do I solve the proportion for the following, 3x-2 0ver 3 = 2x+1 over 4

OpenStudy (phi):

\[ \frac{3x-2}{3}= \frac{2x+1}{4}\]? if so, do you know how to cross multiply?

OpenStudy (anonymous):

no not really

OpenStudy (phi):

ok, do you know how to multiply both sides of the equation by 3?

OpenStudy (anonymous):

3 over 1?

OpenStudy (phi):

yes, but on both sides (to keep things equal) and notice that you can simplify the left side because 3/3 is 1

OpenStudy (anonymous):

ok with you so far

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

i'm sorry I have to get back in a little bit, my son just threw up